of the following,which is greater than 1/2 .2/5,4/7,5/11
Answer by ewatrrr(24785)

 
Hi,
of the following,which is greater than 1/2 = 5/10 = 7/14 = 11/22
2/5 = 4/10
4/7 = 8/14 ****
5/11 = 10/22